OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

how do u

OpenStudy (anonymous):

oh wait i remember

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

nope, where did the x go?

OpenStudy (anonymous):

i thought the x would equal 0 when u solve for x

OpenStudy (anonymous):

i mean when u solve for y

OpenStudy (anonymous):

no, start by subtracting 5x from both sides

OpenStudy (anonymous):

5x -3y=-6

OpenStudy (anonymous):

-3y=-6

OpenStudy (anonymous):

5x -3y=-6 -5x = -5x ----------------- -3y =-6-5x

OpenStudy (anonymous):

now divide both sides by -3

OpenStudy (anonymous):

y=-5

OpenStudy (anonymous):

the equation of a line is given by: y= mx +c where m is the gradient (slope) and C the y -intercept, rearrange your equation in this format and you'll get an answer. so 5x - 3y = -6 becomes -3y = -5x - 6 3y = 5x + 6 y = 5/3 x + 2 so whats your slope?

OpenStudy (anonymous):

-6-5x ------ -3

OpenStudy (anonymous):

-6/-3 - 5x/-3 =2 + 5/3x

OpenStudy (anonymous):

number next to x is slope , other one is y intercept